8-3

Solitaire Companys fixed budget performance report for June follows. The $311,250 budgeted expenses include $249,000 variable expenses and $62,250 fixed expenses. Actual expenses include $71,050 fixed expenses Actual Fixed Budget Results Variances 10,700 $535,000 374,500 $160,500 8,300 Sales (in units) Sales (in dollars) Total expenses Income from operations $415,000 311,250 $103,750 $120,000 F 63,250U 56,750 F Prepare a flexible budget performance report showing any variances between budgeted and actual results. List fixed and variable expenses separately SOLITAIRE COMPANY Flexible Budget Performance Report For Month Ended June 30 Flexible Budget Actual ResultsVariances Fav./Unf Sales 535,000S 535,000 321,000 214,000 o variance avorable avorable nfavorable riable expenses Contribution margin xed expenses 62,250 ncome from operations 151,750 160,500 8,750 Favorable
Answer #1

SOLITAIRE COMPANY

Flexible Budget performance Report

For the month ended June 30

Flexible budget

Actual Result

Variance

Fav./ Unf.

Sales revenue

$   5,35,000.00

$     5,35,000.00

$                 -  

No Variance

Variable Cost

$   3,21,000.00

$     3,03,450.00

$ 17,550.00

Favorable

Contribution margin

$   2,14,000.00

$     2,31,550.00

$ 17,550.00

Favorable

Fixed cost

$      62,250.00

$         71,050.00

$ (8,800.00)

Unfavorable

Operating profit

$   1,51,750.00

$     1,60,500.00

$    8,750.00

Favorable

Working

SOLITAIRE COMPANY

Flexible Budget performance Report

For the month ended June 30

Flexible budget

Actual Result

Variance

Fav./ Unf.

Sales revenue

=415000/8300*10700

535000

535000-535000

No Variance

Variable Cost

=249000/8300*10700

=374500-71050

321000-303450

Favorable